The area in square meters for a triangular
sail is given by A(x) - x2 - 5x - 6.
a) Find A(5).
b) If the height of the sail is x - 3 meters, then what is
the length of the base of the sail?

Answer by RAY100(1637)

A(x) = x^2 -5x -6,,,,,A(5),,just substitute x=5,,,
.
A(5) = (5)^2 -5(5) -6 = 25 -25 -6 =-6,,,possibly should be +6
.
if one side is (x-3),,,,@x=5,,,,side = 5-3 = 2
.
A triangle = (1/2) bh,,,,=(1/2)(2)h= h
.
A=+6 = h
.
Please check original problem and modify as required